OpenStudy (anonymous):

WILL GIVE A MEDAL The graph of y = ax^2 + bx + c passes through the points (-2, 0), (0, -2), and (2, 0). Determine the solution set of 0 = ax^2 + bx + c.

OpenStudy (danjs):

from the point (0,-2) when x=0, y=-2 y = ax^2 + bx + c let x = 0, so y=-2 -2 = 0 + 0 + c c = -2

OpenStudy (danjs):

so far you have y = ax^2 + bx -2

OpenStudy (danjs):

When x = +2 or x=-2, y=0, from the other 2 points

OpenStudy (anonymous):

I thought the solution set was {0}

OpenStudy (jhannybean):

You plug in your points to find your constant value, then use that to find your roots, or x values.

OpenStudy (anonymous):

{-2}, {0}, {-2,2}, and {-2,0,2} are the answer choices

OpenStudy (danjs):

(x-2) (x+2) will be factors of y=0

OpenStudy (anonymous):

So.. what does that mean?

OpenStudy (danjs):

the solution set to 0 = ax^2 + bx + c is where it crosses the x axis,, y = 0

OpenStudy (danjs):

it tells you the points in the problem, (2,0) and (-2,0) those are the x-intercepts, y=0

OpenStudy (anonymous):

So it would be the last one? {-2,0-2}

OpenStudy (anonymous):

*{-2,0,2}

OpenStudy (danjs):

no, not zero, when x=0 y=-2 from the given point (0,-2)

OpenStudy (danjs):

you want the points where y=0, they are the solutions to y = 0 = ax^2 + bx + c

OpenStudy (danjs):

{2 -2}

OpenStudy (anonymous):

Why, thank you so much!

OpenStudy (danjs):

you're welcome, anytime

OpenStudy (danjs):

could find the actual values for a, b, and c, but not needed here, they tell you the x intercepts already

OpenStudy (danjs):

and a polynomial to the 2nd power degree, will have 2 intercepts